Can someone help me solve the radical expression
sqrt3x + 5 = 11
Thank you

Answer by checkley71(8403)

sqrt3x+5=11 or sqrt3x=11-5 opr sqrt3x=6 or squaring both sides we get 3x=36 or x=36/12 or x=12